# Accuracy and Political Bias of News Source Credibility Ratings by Large Language Models

### v2

@@ Abstract
Search engines increasingly leverage large language models (LLMs) to generate direct answers, and AI chatbots now access the Internet for fresh data. As information curators for billions of users, LLMs must assess the accuracy and reliability of different sources. This paper audits eight widely used LLMs from three major providers—OpenAI, Google, and Meta—to evaluate their ability to discern credible and high-quality information sources from low-credibility ones. We find that while LLMs can rate most tested news outlets, larger models more frequently refuse to provide ratings due to insufficient information, whereas smaller models are more prone to hallucination in their ratings. For sources where ratings are provided, LLMs exhibit a high level of agreement among themselves (average Spearman’s $\rho=0.81$), but their ratings align only moderately with human expert evaluations (average $\rho=0.59$). Analyzing news sources with different political leanings in the US, we observe a liberal bias in credibility ratings yielded by all LLMs in default configurations. Additionally, assigning partisan identities to LLMs consistently results in strong politically congruent bias in the ratings. These findings have important implications for the use of LLMs in curating news and political information.
Code & datasetshttps://github.com/osome-iu/llm˙domain˙rating
Introduction
@@
Large language models (LLMs) are being integrated into our information ecosystems, transforming how people seek and consume information. One significant trend is the emergence of AI-augmented search, where an LLM layer is added into traditional search engines to provide direct answers based on search results (Xiong et al. 2024). Major platforms like Google111blog.google/products/search/generative-ai-google-search-may-2024 and Microsoft222blogs.microsoft.com/blog/2023/02/07/reinventing-search-with-a-new-ai-powered-microsoft-bing-and-edge-your-copilot-for-the-web have implemented this feature, and newer products like Perplexity AI and You.com have gained substantial user bases and investments.333techcrunch.com/2024/01/04/ai-powered-search-engine-perplexity-ai-now-valued-at-520m-raises-70m Additionally, AI chatbots are now often connected to the Internet, allowing them to fetch up-to-date information not included in their training data and ground their responses in real-time (Vu et al. 2023). In these systems, LLMs act as information curators, determining what content is shown to billions of users. Recent studies suggest that such integration lowers the information access barrier (Wu et al. 2020) and enables users to perform complex tasks more quickly (Spatharioti et al. 2023; Suri et al. 2024), indicating potential for mainstream adoption.
@@
Despite such advantages, the additional LLM components could introduce new problems (Memon and West 2024) because language models face critical technical challenges, such as hallucinations (Ji et al. 2023), and exhibit various biases (Gallegos et al. 2024). Recent audits of popular AI search engines reveal that their results often contain unsupported claims (Liu, Zhang, and Liang 2023) and biases depending on the queries (Li and Sinnamon 2024). Furthermore, experiments conducted by Sharma, Liao, and Xiao (2024) demonstrate that users may engage with more biased information when interacting with AI search engines, and LLMs with predefined opinions can exacerbate this bias. Beyond these analyses, our understanding of the potential issues of the new LLM layer remains limited.
In this paper we wish to quantify the accuracy with which LLMs can determine the credibility of information sources, a crucial qualification for information curators. To this end, we present experiments that audit eight widely used LLMs from three major providers: OpenAI, Meta, and Google. These models are instructed to provide ratings for over 7,000 websites representing major news sources on the Internet. The accuracy of these ratings is measured by their alignment with human expert evaluations.
@@
After reviewing the related work and describing our methodology, we present the results of our experiments. For most tested news sources, we find that the LLMs can provide ratings as instructed. However, larger models more frequently refuse to rate less popular sources due to insufficient knowledge, while smaller models are more prone to hallucinating responses. The LLM ratings show a high level of agreement among themselves, even though they are trained by different providers. However, their ratings only moderately correlate with those of human experts. Focusing on news sources with clear political leanings in the US, we find that assigning partisan identities to LLMs consistently biases their ratings in favor of sources with congruent leanings. Additionally, LLMs exhibit a liberal bias in their default configurations.
@@
Our findings suggest that while LLMs have the potential to judge the credibility of information sources, state-of-the-art models from different companies exhibit common problems. Their lack of knowledge about unpopular information sources presents challenges in dealing with data voids (Boyd and Golebiewski 2018; Aslett et al. 2024). Furthermore, inaccurate LLM ratings due to issues like hallucinations and biases could inadvertently amplify misinformation while suppressing high-credibility sources. When handling political information, the inherent partisan biases in LLMs could exacerbate echo chambers and polarization (Cinelli et al. 2021; Conover et al. 2021). Therefore, we caution against relying solely on LLMs for information curation and call for further evaluations and improvements to ensure their reliability and accuracy.
Related Work
@@
Our study follows a long line of scholarship aimed at understanding how online platforms affect user information diets (Rahwan et al. 2019) and their contributions to issues such as the proliferation of misinformation (Lazer et al. 2018), echo chambers (Cinelli et al. 2021), and polarization (Conover et al. 2021). As prominent platforms for information-seeking, search engines have been scrutinized by many researchers evaluating if their ranking algorithms favor certain news sources (Robertson et al. 2018; Trielli and Diakopoulos 2019). Social media platforms are another focus of attention. Previous studies address topics such as political biases and the amplification of malicious content in various platforms, including Twitter/X (Chen et al. 2021), Facebook (Bakshy, Messing, and Adamic 2015), and YouTube (Hosseinmardi et al. 2021).
@@
The emergence of LLMs such as ChatGPT has initiated a new line of research auditing these models to ensure they are deployed in ways that are ethical, safe, and accountable (Mökander et al. 2023). Recent studies show that LLMs can perpetuate discrimination due to their inherent biases (Bai et al. 2024; Salinas et al. 2023) and stereotypes (Abid, Farooqi, and Zou 2021; Cheng, Durmus, and Jurafsky 2023). In addition, LLMs have often been found to exhibit a liberal bias in the political context (Hartmann, Schwenzow, and Witte 2023; Santurkar et al. 2023).
@@
As LLMs are increasingly integrated into online platforms, there is a growing need for research to evaluate the potential risks associated with this use. Several recent studies have begun to address this topic by auditing AI search engines such as Bing Chat, Google Bard, and Perplexity AI. By querying these tools with various search phrases, Liu, Zhang, and Liang (2023) show that the responses often contain unsupported claims (Liu, Zhang, and Liang 2023); Li and Sinnamon (2024) identify sentiment and geographic biases; and Urman and Makhortykh (2023) reveal significant disparities across different services when dealing with political information. Through experiments in which participants interact with AI search engines in a lab environment, Sharma, Liao, and Xiao (2024) find that users tend to engage with more biased information, and opinionated LLMs can exacerbate this bias. Despite these research efforts, our understanding of the impact of LLMs as information curators remains limited.
Methods
Model Selection
@@ Methods
We audit eight state-of-the-art models from three major providers. These models have been deployed through various services and systems, which directly or indirectly interact with billions of users across the world every day.
@@
For OpenAI, we test GPT-4o mini (gpt-4o-mini-2024-07-18), GPT-4o (gpt-4o-2024-05-13), and GPT-4 (gpt-4-turbo-2024-04-09). Although OpenAI has not released the number of parameters used by these models, the price and response time suggest that GPT-4o mini is the smallest and GPT-4 is the largest. These models are available to users through ChatGPT and serve as the backend for AI search engines like Bing. In our experiment, we query OpenAI’s API endpoints directly.
@@
For Meta, we test Llama 3.1 models with 8B, 70B, and 405B parameters (Llama Team, AI at Meta 2024). Meta has incorporated these models into its products, such as WhatsApp and Facebook, allowing users to interact with them directly. Due to their open-weight nature, Llama models have also been adopted in many third-party services and applications. Here, we query the Llama 3.1 models hosted by together.ai through its API endpoints.
@@
System prompt You are an assistant to determine the credibility of websites.
Then we provide the model with the task instructions:
@@
Task instruction Rate the website’s credibility: {domain}, on a scale between 0 and 1, where 0 means very low credibility and 1 means very high credibility. The assistant returns a rating of $-1$ when the assistant has no knowledge of the website; otherwise, the assistant should provide the best estimation. In addition to the rating, the assistant should also provide a short explanation.
@@
In each query, {domain} is replaced with the news website of interest. Note that the complete instructions also include specifications for the output; see the Appendix for details.
When evaluating the impact of assigning political identities to LLMs, we append the following instruction to the system prompt:
Political identity assignment You identify as {identity} on the US political spectrum.
@@
In the experiments, {identity} is replaced with the one of “a Democrat,” “an Independent,” or “a Republican” to induce personas with different political views.
News Outlet Credibility Ratings
@@
In this study, we adopt ratings of news source credibility compiled by Lin et al. (2023). The authors analyze the news ratings from six sources, including NewsGuard,444newsguardtech.com/ratings/rating-process-criteria adfontesmedia.com, Iffy index of unreliable sources,555iffy.news MBFC,666mediabiasfactcheck.com/methodology and two lists compiled by professional fact-checkers and researchers (Pennycook and Rand 2019; Lasser et al. 2022). The comparison of these ratings reveals a high level of consistency. Using an ensemble method, they generate an aggregate list that contains credibility ratings for 11,520 websites.
Website Popularity Ranking
@@
We adopt the Tranco list to measure website popularity (Pochat et al. 2018). This list combines the website ranking information from multiple sources, including Alexa777alexa.com/topsites and Cisco Umbrella.888umbrella-static.s3-us-west-1.amazonaws.com/index.html The Tranco list is updated on a routine basis, and for this study we use the snapshot from September 2022. This snapshot contains the ranks of the top 5.5 million websites worldwide. Websites not on the list typically have limited viewership or have been deactivated.
Merging the Tranco list with the list from Lin et al. (2023), we end up with 7,523 websites with credibility ratings and popularity ranks. In the following, we refer to these as the “human expert ratings” and use them as a reference to measure the accuracy of LLM ratings.
Source Political Bias Score
@@
To quantify partisan bias of the news sources in the US political context, we leverage the scores generated by Robertson et al. (2018). The authors first construct a panel of over half a million Twitter accounts matched with US voter registration records and then measure the partisan bias of the audience of each source. The results are scores between $-1$ and $+1$, where a score of $-1$ means that a source is shared exclusively by Democrats, and a score of $+1$ means that a source is shared exclusively by Republicans. The authors show that their scores correlate highly with other ratings, such as the ones produced by Bakshy, Messing, and Adamic (2015), Budak, Goel, and Rao (2016), and allsides.com. Merging the list from Robertson et al. (2018) with the human expert ratings of credibility, we obtain 2,683 sources with political leanings to analyze the political biases of LLMs.
Results
LLM Response Analysis
@@ Results
As described in the Methods section, we first run the 7,523 news sources through all eight LLMs using the same prompt in the default configuration (no assigned political identities). In most cases, the LLMs are able to respond with the required content in the specified format. When errors occur, we rerun the queries until we obtain the desired outputs.
Take reuters.com as an example. GPT-4o provides the following response:
GPT-4o’s assessment of reuters.com Rating: 1.0 Explanation: Reuters is a well-established international news organization known for its accurate and unbiased reporting. It is widely respected in the journalism community and has a long history of providing reliable news.
@@
All other models give Reuters credibility scores over 0.9 with similar explanations (complete responses available in the Appendix). These responses indicate that LLMs recognize news outlets from their websites, possess information about them, and can provide credibility ratings.
@@
Figure 1: Relationship between source popularity and the responses of LLMs. The left axes and the lines show the percentages of sources for which each LLM provides ratings. The dashed lines indicate the overall percentages, whereas the solid lines illustrate the results for different source ranking deciles. The right axes and the dots represent the Spearman correlation coefficients between LLM ratings and human expert ratings in different source ranking deciles. Sources in larger ranking deciles are less popular.
@@
When the LLMs lack sufficient information about a given source, they respond with a rating of $-1$, as instructed. In Figure 1, we illustrate the percentage of sources for which each LLM provides ratings (dashed lines). We find that within each family, larger models are more likely to indicate insufficient information about the given sources and to refuse to rate them. We hypothesize that LLMs tend to lack knowledge about unpopular news sources. To confirm this, we split the sources into 10 deciles based on the popularity ranking and calculate the percentage of sources with LLM ratings within each decile. The solid lines in Figure 1 support our hypothesis for all models.
@@
Figure 2: Percentage of responses based on hallucinations among 200 manually annotated cases for each LLM.
@@
Figure 1 also shows that Llama models can provide ratings for more sources compared to GPT and Gemini models. However, LLMs are known to suffer from hallucinations, where they generate baseless responses to user requests (Ji et al. 2023). To gain deeper insights into the ratings generated by these LLMs, we randomly select 200 sources and manually analyze the responses from all eight models. Through this process, we identify two common types of hallucinations. First, the LLMs can mistakenly associate a website with the wrong organization. For example, five out of eight LLMs identify aldf.com, the official website of the “American Lyme Disease Foundation,” as belonging to the “Animal Legal Defense Fund” (aldf.org). The complete responses of the LLMs can be found in the Appendix. Second, the LLMs can fail to identify the websites yet still generate ratings for them.
@@
In Figure 2, we illustrate the percentage of unambiguous hallucinations among the annotated cases for each LLM. The results show that smaller models tend to hallucinate more frequently within each family. Across different providers, Llama and Gemini models exhibit higher rates of hallucination. It is important to note that even when the models are not hallucinating, they may still provide inaccurate ratings for the sources due to other issues. Next, we assess the accuracy of the LLM ratings by comparing them with those given by human experts.
Rating Accuracy

### v3

@@ Abstract.
Search engines increasingly leverage large language models (LLMs) to generate direct answers, and AI chatbots now access the Internet for fresh data. As information curators for billions of users, LLMs must assess the accuracy and reliability of different sources. This paper audits nine widely used LLMs from three leading providers—OpenAI, Google, and Meta—to evaluate their ability to discern credible and high-quality information sources from low-credibility ones. We find that while LLMs can rate most tested news outlets, larger models more frequently refuse to provide ratings due to insufficient information, whereas smaller models are more prone to making errors in their ratings. For sources where ratings are provided, LLMs exhibit a high level of agreement among themselves (average Spearman’s $\rho=0.79$), but their ratings align only moderately with human expert evaluations (average $\rho=0.50$). Analyzing news sources with different political leanings in the US, we observe a liberal bias in credibility ratings yielded by all LLMs in default configurations. Additionally, assigning partisan roles to LLMs consistently induces strong politically congruent bias in their ratings. These findings have important implications for the use of LLMs in curating news and political information.
Large language models, AI search engines, news credibility, political bias
copyright: none
1. Introduction
@@
Large language models (LLMs) are being integrated into our information ecosystems, transforming how people seek and consume information. One significant trend is the emergence of AI-augmented search, where an LLM layer is added to traditional search engines to provide direct answers based on search results (Xiong et al., 2024). Major platforms like Google111blog.google/products/search/generative-ai-google-search-may-2024 and Microsoft222blogs.microsoft.com/blog/2023/02/07/reinventing-search-with-a-new-ai-powered-microsoft-bing-and-edge-your-copilot-for-the-web have implemented this feature, and newer products like Perplexity AI and You.com have gained substantial user bases and investments.333techcrunch.com/2024/01/04/ai-powered-search-engine-perplexity-ai-now-valued-at-520m-raises-70m Additionally, AI chatbots are now often connected to the Internet, allowing them to fetch up-to-date information not included in their training data and ground their responses in real-time (Vu et al., 2023). In these systems, LLMs act as information curators, determining what content is shown to billions of users. Recent studies suggest that such integration lowers the information access barrier (Wu et al., 2020) and enables users to perform complex tasks more quickly (Spatharioti et al., 2023; Suri et al., 2024), indicating potential for mainstream adoption.
@@
Despite such advantages, the additional LLM components could introduce new problems (Memon and West, 2024) because language models face critical technical challenges, such as hallucinations (Ji et al., 2023), and exhibit various biases (Gallegos et al., 2024). Recent audits of popular AI search engines reveal that their results often contain unsupported claims (Liu et al., 2023) and biases depending on the queries (Li and Sinnamon, 2024). Furthermore, experiments conducted by Sharma et al. (2024) demonstrate that users may engage with more biased information when interacting with AI search engines, and LLMs with predefined opinions can exacerbate this bias. Beyond these analyses, our understanding of the potential issues of the new LLM layer remains limited.
This study investigates whether LLMs exhibit errors and political biases when evaluating the credibility of information sources, as these issues could compromise their effectiveness as information curators. We conduct comprehensive experiments to audit nine prominent LLMs from three leading providers: OpenAI, Meta, and Google. Our methodology involves instructing these models to evaluate and rate over 7,000 websites that represent major news sources across the Internet. We assess the accuracy of these LLM-generated ratings by comparing them with evaluations provided by human experts.
@@
After reviewing the related work and describing our methodology, we present the results of our experiments. For most tested news sources, we find that the LLMs can provide ratings as instructed. However, larger models more frequently refuse to rate less popular sources due to insufficient knowledge, while smaller models are more prone to make errors. The LLM ratings show a high level of agreement among themselves, even though they are trained by different providers. However, their ratings only moderately correlate with those of human experts. Focusing on news sources with clear political leanings in the US, we find that assigning partisan roles (e.g., Democrat and Republican) to LLMs consistently biases their ratings in favor of sources with congruent leanings. Additionally, LLMs exhibit a liberal bias in their default configurations.
@@
Our findings suggest that while LLMs have the potential to judge the credibility of information sources, state-of-the-art models from different companies exhibit common problems. Their lack of knowledge about unpopular information sources presents challenges in dealing with data voids (Boyd and Golebiewski, 2018; Aslett et al., 2024). Furthermore, inaccurate LLM ratings due to issues like errors and biases could inadvertently amplify misinformation while suppressing high-credibility sources. When handling political information, the inherent partisan biases in LLMs could exacerbate echo chambers and polarization (Cinelli et al., 2021; Conover et al., 2021). Therefore, we caution against relying solely on LLMs for information curation and call for further evaluations and improvements to ensure their reliability and accuracy.
2. Related Work
@@
Our study follows a long line of scholarship aimed at understanding how online platforms affect user information diets (Rahwan et al., 2019) and their contributions to issues such as the proliferation of misinformation (Lazer et al., 2018), echo chambers (Cinelli et al., 2021), and polarization (Conover et al., 2021). As prominent platforms for information-seeking, search engines have been scrutinized by many researchers evaluating if their ranking algorithms favor certain news sources (Robertson et al., 2018; Trielli and Diakopoulos, 2019). Social media platforms are another focus of attention. Previous studies address topics such as political biases and the amplification of malicious content in various platforms, including Twitter/X (Chen et al., 2021), Facebook (Bakshy et al., 2015), and YouTube (Hosseinmardi et al., 2021).
@@
The emergence of LLMs such as ChatGPT has initiated a new line of research auditing these models to ensure they are deployed in ways that are ethical, safe, and accountable (Mökander et al., 2023). Recent studies show that LLMs can perpetuate discrimination due to their inherent biases (Bai et al., 2024; Salinas et al., 2023) and stereotypes (Abid et al., 2021; Cheng et al., 2023). In addition, LLMs have often been found to exhibit a liberal bias in the political context (Hartmann et al., 2023; Santurkar et al., 2023).
@@
As LLMs are increasingly integrated into online platforms, there is a growing need for research to evaluate the potential risks associated with this use. Several recent studies have begun to address this topic by auditing AI search engines such as Bing Chat, Google Bard, and Perplexity AI. By querying these tools with various search phrases, Liu et al. (2023) show that the responses often contain unsupported claims (Liu et al., 2023); Li and Sinnamon (2024) identify sentiment and geographic biases, and Urman and Makhortykh (2025) reveal significant disparities across different services when dealing with political information. Through experiments in which participants interact with AI search engines in a lab environment, Sharma et al. (2024) find that users tend to engage with more biased information, and opinionated LLMs can exacerbate this bias. Despite these research efforts, our understanding of the impact of LLMs as information curators remains limited.
Our study evaluates the capability of LLMs to assess the credibility of information sources. Traditionally, this assessment has been conducted by human experts. Organizations such as NewsGuard444newsguardtech.com/ratings/rating-process-criteria and MBFC555mediabiasfactcheck.com/methodology employ teams of professional fact-checkers and media literacy experts who systematically evaluate and rate the credibility of diverse information sources. Recent research has started to explore automated approaches to this challenge. Carragher et al. (2024) investigates the application of web graphs for automated detection of unreliable domains, while Williams et al. (2024) demonstrates enhanced performance by combining signals from web graphs and social media. Unlike these studies, we instruct LLMs to finish the tasks in a zero-shot manner in our experiments without training or providing them with additional data.
3. Methods
3.1. Model Selection
@@
We audit nine state-of-the-art models from three leading providers. These models have been deployed through various services and systems, which directly or indirectly interact with billions of users across the world every day.
@@
For OpenAI, we test GPT-4o mini (gpt-4o-mini-2024-07-18), GPT-4o (gpt-4o-2024-05-13), and GPT-4 (gpt-4-turbo-2024-04-09). Although OpenAI has not released the number of parameters used by these models, the price and response time suggest that GPT-4o mini is the smallest and GPT-4 is the largest. Additionally, we test the o1-mini (o1-mini-2024-09-12) model, which is designed for complex reasoning tasks by explicitly incorporating chain-of-thought into its inference process (OpenAI, 2024). Such a design improves o1-mini’s performance on certain tasks, but its number of parameters is not comparable to the other models. These OpenAI models are available to users through ChatGPT and serve as the backend for AI search engines like Bing. In our experiment, we query OpenAI’s API endpoints directly.
@@
For Meta, we test Llama 3.1 models with 8B, 70B, and 405B parameters (Llama Team, AI at Meta, 2024). Meta has incorporated these models into its products, such as WhatsApp and Facebook, allowing users to interact with them directly. Due to their open-weight nature, Llama models have also been adopted in many third-party services and applications. Here, we query the Llama 3.1 models hosted by together.ai through its API endpoints.
@@ 3.2. Prompts
System prompt:
You are an assistant to determine the credibility of websites.
Then, we provide the model with the task instructions:
Task instruction:
@@
Rate the website’s credibility: {domain}, on a scale between 0 and 1, where 0 means very low credibility and 1 means very high credibility. The assistant returns a rating of $-1$ when the assistant has no knowledge of the website; otherwise, the assistant should provide the best estimation. In addition to the rating, the assistant should also provide a short explanation.
@@
In each query, {domain} is replaced with the news website of interest. Here, we only provide the LLMs with the domain names to test if the LLMs have knowledge of the news sources encoded. Note that the complete instructions also include specifications for the output; see the Appendix for details.
The prompt above tests the default configuration of LLMs. To further investigate how easily the responses of LLMs can be steered, we also test the impact of assigning partisan roles to LLMs in the prompts. Specifically, we append the following instructions to the system prompt:
Partisan role assignment:
You identify as {role} on the US political spectrum.
@@
In the experiments, {role} is replaced with the one of “a Democrat,” “an Independent,” or “a Republican” to induce different political biases.
3.3. News Outlet Credibility Ratings
@@
In this study, we adopt ratings of news source credibility compiled by Lin et al. (2023). The authors analyze the news ratings from six sources, including NewsGuard,666newsguardtech.com/ratings/rating-process-criteria adfontesmedia.com, Iffy index of unreliable sources,777iffy.news MBFC,888mediabiasfactcheck.com/methodology and two lists compiled by professional fact-checkers and researchers (Pennycook and Rand, 2019; Lasser et al., 2022). The comparison of these ratings reveals a high level of consistency. Using an ensemble method, they generate an aggregate list that contains credibility ratings for 11,520 websites.
3.4. Website Popularity Ranking
@@
Our examination of the news source list from Lin et al. (2023) reveals that it contains websites that are no longer active so testing with them is not meaningful. To remove these websites, we leverage the Tranco list that measures website popularity (Le Pochat et al., 2019). The Tranco list combines the website ranking information from multiple sources, including Alexa999alexa.com/topsites and Cisco Umbrella.101010umbrella-static.s3-us-west-1.amazonaws.com/index.html It is updated on a routine basis, and for this study, we use the snapshot from September 2022. This snapshot contains the ranks of the top 5.5 million websites worldwide. Its intersection with the list from Lin et al. (2023) contains 7,523 websites. In the following, we refer to these as the “human expert ratings” and use them as a reference to measure the accuracy of LLM ratings.
3.5. Source Political Bias Score
@@
To quantify partisan bias of the news sources in the US political context, we leverage the scores generated by Robertson et al. (2018). The authors first construct a panel of over half a million Twitter accounts matched with US voter registration records and then measure the partisan bias of the audience of each source. The results are scores between $-1$ and $+1$, where a score of $-1$ means that a source is shared exclusively by Democrats, and a score of $+1$ means that a source is shared exclusively by Republicans. The authors show that their scores correlate highly with other ratings, such as the ones produced by Bakshy et al. (2015), Budak et al. (2016), and allsides.com. Merging the list from Robertson et al. (2018) with the human expert ratings of credibility, we obtain 2,683 sources with political leanings to analyze the political biases of LLMs.
4. Results
4.1. LLM Response Analysis
@@
As described in the Methods section, we first run the 7,523 news sources through all nine LLMs using the same prompt in the default configuration (no partisan roles). In most cases, the LLMs respond with the required content in the specified format. When errors occur, we rerun the queries until we obtain the desired outputs. The code and data are available at our GitHub repository111111github.com/osome-iu/llm_domain_rating.
@@
Take reuters.com as an example. GPT-4o provides the following response: “Rating: 1.0; Explanation: Reuters is a well-established international news organization known for its accurate and unbiased reporting. It is widely respected in the journalism community and has a long history of providing reliable news.” All other models give Reuters credibility scores over 0.9 with similar explanations (complete responses available in the Appendix). These responses indicate that LLMs recognize news outlets from their websites, possess information about them, and can provide credibility ratings.
@@
Figure 1. Relationship between source popularity and the responses of LLMs. The left axes and the lines show the percentages of sources for which each LLM provides ratings. The dashed lines indicate the overall percentages, whereas the solid lines illustrate the results for different source ranking deciles. The right axes and the dots represent the Spearman correlation coefficients between LLM ratings and human expert ratings in different source ranking deciles. Sources in larger ranking deciles are less popular.
[…]
Relationship between source popularity and the responses of LLMs.
@@
When the LLMs lack sufficient information about a given source, they respond with a rating of $-1$, as instructed. In Figure 1, we illustrate the percentage of sources for which each LLM provides ratings (dashed lines). We find that within each family, larger models are more likely to refuse to rate sources due to insufficient information. We hypothesize that LLMs tend to lack knowledge about unpopular news sources. To confirm this, we split the sources into 10 deciles based on the popularity ranking and calculate the percentage of sources with LLM ratings within each decile. The solid lines in Figure 1 support our hypothesis for all models.
@@
Figure 2. Percentage of errors among 200 manually annotated cases for each LLM.
[…]
Percentage of errors among 200 manually annotated cases for each LLM.
@@
Figure 1 also shows that Llama models can provide ratings for more sources compared to OpenAI and Gemini models. However, LLMs are known to suffer from hallucinations, where they generate baseless responses to user requests (Ji et al., 2023). To gain deeper insights into the ratings generated by these LLMs, we randomly select 200 sources and manually analyze the responses from all nine models. Through this process, we identify two common types of unambiguous errors. First, the LLMs can mistakenly associate a website with the wrong organization. For example, five out of nine LLMs identify aldf.com, the official website of the “American Lyme Disease Foundation,” as belonging to the “Animal Legal Defense Fund” (aldf.org). The complete responses of the LLMs can be found in the Appendix. Second, an LLM can fail to identify a website and still generate a rating for it.
@@
In Figure 2, we illustrate the percentage of unambiguous errors among the annotated cases for each LLM. The results show that smaller models tend to make more errors within each family. In the OpenAI family, o1-mini exhibits the highest rate of errors among all models. However, as mentioned in the Methods section, its number of parameters is not comparable to the other models due to its special design. Across different providers, Llama and Gemini models exhibit higher rates of errors in general. It is important to note that even when the models correctly recognize the sources, they may still provide inaccurate ratings due to other issues. Next, we assess the accuracy of the LLM ratings by comparing them with those given by human experts.
4.2. Rating Accuracy